# On an approach to canonicalizing elliptic Feynman integrals

### v1

@@ Acknowledgements
We would like to thank Ming Lian for collaborations in the early stage of this work. This work was supported in part by the National Natural Science Foundation of China under Grant No. 12375097, 12347103, and the Fundamental Research Funds for the Central Universities.
@@ Appendix A Construct $\varepsilon$-factorized basis with constant period matrix
$\phi_{i-3}\to\big{(}\,\mathcal{T}_{1}.\vec{\phi}\,\big{)}_{i-3}=\phi_{i-3}+\left[F(\sin^{-1}a_{i},\lambda)-E(\sin^{-1}a_{i},\lambda)\right]\phi_{1}-F(\sin^{-1}a_{i},\lambda)\phi_{n-1}\,,$ (105)
for $15\leq i\leq n+$ 1, with $a_{i}=\sqrt{e_{i}/\lambda}$, and we have implicitly introduced the rotation matrix $\mathcal{T}_{1}$ acting on the basis. The transformation carries over to the integral level, namely, $I_{i-3}\to\big{(}\,\mathcal{T}_{1}.\vec{I}\,\big{)}_{i-3}$, and we will suppress it for brevity. After such a transformation, it is straightforward to verify that the derivatives (with respect to $\lambda$ or $e_{i}$) of master integrals $I_{j}$’s with $23\leq j\leq n-$ 2, are $\varepsilon$-factorized under the maximal cut.
@@ Appendix B Canonical basis with Jacobi normal form
$u_{J}(t)=P_{4J}(t)^{-1/2}\prod_{i=1}^{n}(t-a_{i})^{-\beta_{i}\varepsilon}\,,$ (111)

### v2

@@ Acknowledgements
We would like to thank Ming Lian for collaborations in the early stage of this work. This work was supported in part by the National Natural Science Foundation of China under Grant No. 12375097, 12347103, Science Foundation of China University of Petroleum, Beijing (No.2462025YJRC019), and the Fundamental Research Funds for the Central Universities.
@@ Appendix A Construct $\varepsilon$-factorized basis with constant period matrix
$\phi_{i-3}\to\big{(}\,\mathcal{T}_{1}\cdot\vec{\phi}\,\big{)}_{i-3}=\phi_{i-3}+\left[F(\sin^{-1}a_{i},\lambda)-E(\sin^{-1}a_{i},\lambda)\right]\phi_{1}-F(\sin^{-1}a_{i},\lambda)\phi_{n-1}\,,$ (105)
for $15\leq i\leq n+$ 1, with $a_{i}=\sqrt{e_{i}/\lambda}$, and we have implicitly introduced the rotation matrix $\mathcal{T}_{1}$ acting on the basis. The transformation carries over to the integral level, namely, $I_{i-3}\to\big{(}\,\mathcal{T}_{1}\cdot\vec{I}\,\big{)}_{i-3}$, and we will suppress it for brevity. After such a transformation, it is straightforward to verify that the derivatives (with respect to $\lambda$ or $e_{i}$) of master integrals $I_{j}$’s with $23\leq j\leq n-$ 2, are $\varepsilon$-factorized under the maximal cut.
@@ Appendix B Canonical basis with Jacobi normal form
$u_{J}(t)=[P_{4J}(t)]^{-1/2}\prod_{i=1}^{n}(t-a_{i})^{-\beta_{i}\varepsilon}\,,$ (111)